OpenStudy (anonymous):

Which system of equations is (–3, 19) and (4, –23) equal to?

OpenStudy (anonymous):

A.y=x^2+7x+11 y=6x-1 B. y=x^2-7x+11 y=6x+1 C.-x^2+7x-11 y=6x-1 D.y=x^2-7x-11 y=-6x+1

OpenStudy (whpalmer4):

You mean which system of equations have those two pairs of numbers as solutions. Plug each pair of numbers into each of the equations in a given system. If they all produce true number sentences, then it is a solution; otherwise, it isn't. Given that you have a multiple choice format answer here, you might try plugging the numbers into the second equation in each one first, so you can discard candidates more easily.

OpenStudy (anonymous):

that was what i needed to know how to do it thank u

OpenStudy (whpalmer4):

What did you get for the answer?

OpenStudy (anonymous):

i believe it is c

OpenStudy (whpalmer4):

No....you couldn't possibly have tried the points in the equations successfully and chosen C, I'm afraid :-( C has two equations:\[y=-x^2+7x-11\]and\[y=6x-1\] We have two points to try: \((-3,19) \text{ and } (4,-23)\) \[y = -x^2+7x-11\]\[19 = -(-3)^2+7(-3)-11\]\[19 = -9 -21 -11\]Uh, no. \[y = 6x-1\]\[19=6(-3)-1\]\[19=-19\]Also no. You're trying to find a set of equations that make valid number sentences when you plug in the point. Maybe you tried the other point? \[-23 = -(4)^2+7(4)-11\]\[-23=-16+28-11\]\[-23 = 1\]Nope, doesn't look like you tried the other point, either. \[-23 = 6(4) -1\]\[-23 = 23\]